# Learning Representations of Event Time Series with Sparse Autoencoders for Anomaly Detection, Similarity Search, and Unsupervised Classification

### v1

@@ 2.1 Tensor Representation of Event Time Series
Assume that each event time series consists of a sequence of timestamps $\boldsymbol{t}=\{t_{k}\}_{k=1}^{N}$ and associated event modality values $\boldsymbol{E}=\{E_{k}\}_{k=1}^{N}$. In the context of X-ray astronomy, $\boldsymbol{t}$ would correspond to the photon arrival times and $\boldsymbol{E}$ to the photon energies. The total duration of the event time series is given by $T=t_{N}-t_{1}$, and the event values lie in the range $[E_{\min},E_{\max}]$, where $E_{\min}=\min_{k}E_{k}$ and $E_{\max}=\max_{k}E_{k}$. To standardize the inputs across event time series, we normalize the time axis as
@@
These three normalized axes—event time ($\boldsymbol{\tau}$), transformed modality ($\boldsymbol{\epsilon}$), and inter-event interval ($\boldsymbol{\Delta\tau}$)—are discretized into fixed-width bins with $n_{\tau}$, $n_{\epsilon}$, and $n_{\Delta\tau}$ bins, respectively. Each event is assigned to a voxel in the resulting three-dimensional histogram based on its coordinates in this normalized space. The number of bins along each axis can be chosen flexibly to achieve the desired level of temporal or modality-specific resolution, depending on the application, domain, and characteristics of the data. This allows the representation to be tailored to the appropriate scale and granularity for a given task, while still providing a consistent input format for models, even when the dataset contains sequences of variable length and irregular temporal structure. This results in a tensor $\boldsymbol{X}_{i}\in\mathbb{R}^{n_{\tau}\times n_{\epsilon}\times n_{\Delta\tau}}$, which provides a standardized, structured input format for each of the original event time series $\boldsymbol{x}_{i}$. Each tensor is then collected into a final processed dataset $\{\boldsymbol{X}_{i}\}^{n}_{i=1}$. We also refer to these tensors as $E$$t$$dt$ cubes. If we were to only bin along time $\boldsymbol{\tau}$ and modality $\boldsymbol{\epsilon}$, we would obtain two-dimensional tensors $\boldsymbol{X}_{i}\in\mathbb{R}^{n_{\tau}\times n_{\epsilon}}$, which we refer to as $E$$t$ maps. Appendix A shows $E$$t$ maps and $E$$t$$dt$ cubes for an X-ray flare, dip and pulsating source.
Figure 1: Two-dimensional t-SNE projection of the learned latent space from the SAE applied on the $E$$t$-$dt$ cubes. Panel A: Points are color-coded by the variability index of the corresponding X-ray sources. Panel B: Points are color-coded by the hard-to-soft X-ray hardness ratio. Panel C: Known dips, flares, and pulsating sources (crosses) occupy distinct clusters in the embedding space, enabling the identification of new transient candidates via clustering and similarity searches.
@@ 2.2 Sparse Representation Learning
We now aim to learn compact and informative representations from the $E$$t$$dt$ cubes. To this end, we employ a SAE—an unsupervised neural network model designed to extract structured, low-dimensional features from high-dimensional inputs. By learning to reconstruct the input data while enforcing sparsity in the latent space, the SAE identifies salient patterns and compresses each tensor into a representation that preserves essential temporal and modality-specific features, while remaining insensitive to irrelevant variation (Ranzato et al., 2007; Ng et al., 2011). This includes robustness to noise, systematics, observational artifacts, and context-dependent features introduced by the data acquisition process. Instead, the learned representations are encouraged to capture the most important latent factors that reflect the underlying physical characteristics and meaningful signals encoded in the time series, rather than superficial or incidental aspects of the data.
@@
where $\lambda>0$ controls the strength of the sparsity regularization. Once trained, we discard the decoder and use the encoder $\phi_{\theta}$ as a standalone feature extractor. The resulting latent vectors $\boldsymbol{z}_{i}$ serve as learned representations of the original event time series and can be directly used for downstream tasks such as clustering, anomaly detection, or similarity-based retrieval. A similar procedure is followed when using the $E$$t$ maps as inputs to a SAE.
@@ 3 Experiments on Downstream Tasks
We apply our method to data from the Chandra Source Catalog (CSC) version 2.1 (Evans et al., 2024). Specifically, we use region event files, which list individual photon events for a specific source in a given observation—each with associated arrival time and energy. The dataset comprises 95,473 event files from 58,932 unique X-ray sources. For further details on the dataset, please refer to Dillmann et al. 2025. Using the terminology introduced in Section 2.1, the photon arrival times correspond to $\boldsymbol{t}$ and the photon energies to $\boldsymbol{E}$. We choose the event transformation function $f(\cdot)$ to be the base-10 logarithm, i.e., $f(\boldsymbol{E})=\log_{10}(\boldsymbol{E})$, which is commonly used in X-ray astronomy. For the $E$$t$$dt$ cubes, we choose a dimensionality of $(n_{\tau},n_{\epsilon},n_{\Delta\tau})=(24,16,16)$. For the SAE, we choose a sparsity strength of $\lambda=0.1$ and its latent vector dimension is 24. We also run the experiments described in this section on the $E$$t$ maps with a dimensionality of $(n_{\tau},n_{\epsilon})=(24,16)$. In this case, we use a SAE with convolution layers and a latent vector dimension of 12. More details are provided in Appendix B.
@@
We train a vanilla XGBoost classifier and regressor (Chen & Guestrin, 2016) on the learned latent representations to distinguish highly variable X-ray sources (variability index $>$ 6) from non-variable ones (variability index $\leq$ 6), and to predict the hardness ratio of each source. The performances are summarized in Table 1. The classifier achieves a high variability classification accuracy of 0.97, outperforming the method proposed in Song et al. (2025), and the hardness ratio regressor attains a slightly reduced R2 score of 0.76, demonstrating that the learned representations encode temporal variability well and retain substantial predictive power for spectral hardness.
@@ 4 Conclusion
We introduce a unified framework for learning sparse representations from event time series to enable a variety of different downstream tasks. We represent irregular event time series as $E$$t$$dt$ cube tensors and employ a SAE to learn sparse embeddings that capture physically meaningful variation while increasing their robustness to noise and other irrelevant features. The method requires minimal manual tuning and scales to large datasets across various domains, and serves as a foundation for subsequent analysis like anomaly detection, similarity search, and unsupervised classification. Applied to archival Chandra X-ray data, the learned representations encode temporal and spectral information and enabled the discovery of a new extragalactic FXT and a TDE from a hyperluminous supersoft X-ray source, both previously missed by other transient search methods. Future work involves optimizing the resolution of the $E$$t$$dt$ cubes, the deployment of more expressive SAE architectures and the application of the method to a variety of other scientific datasets.
@@ Acknowledgements
Appendix A $E$-$t$ Maps and $E$-$t$-$dt$ Cubes Visualizations
Figure 3 displays the light curves, $E$$t$ maps, and $E$$t$$dt$ cubes for event files featuring a dip, a flare, and pulsations, respectively.
Figure 3: The top row shows 300 s bin light curves from event files featuring a dip (blue), a flare (red), and pulsations (green), respectively. The middle row shows the corresponding $E$$t$ maps, and the bottom row shows the corresponding $E$$t$$dt$ cubes for these event files.
@@ Appendix B SAE Architecture and Training
The architecture of the experimental SAE for the $E$$t$$dt$ cube inputs is summarized in Table 2. For the experiments presented in this paper, we flatten the input and use fully-connected layers for computational reasons. We recommend experimenting with more sophisticated SAE architectures to fully leverage the value of the $E$$t$$dt$ cubes and the presented SAE representation learning approach. Starting with the original dataset of 95,473 samples, we first apply a 90/10 split to obtain 85,925 samples for training and validation, and 9,548 for testing. We then further split the training and validation subset using an 80/20 ratio, resulting in 68,740 training samples and 17,185 validation samples. The model is trained for up to 200 epochs with a batch size of 1024. Optimization is performed using Adam (Kingma & Ba, 2014) with an initial learning rate of $0.01$. A learning rate scheduler reduces the rate by a factor of 10 if the validation loss plateaus for more than 10 epochs. We also apply early stopping based on the validation loss, which halts training if no improvement is observed for 25 consecutive epochs and restores the best-performing model weights. A similar procedure is followed for the experimental SAE with convolution layers summarized in Table 3 that we used for the $E$$t$ map inputs.
Table 2: Summary of the encoder architecture of the fully-connected autoencoder used to extract informative features from the $E$-$t$-$dt$ cubes. Each layer uses Leaky ReLU activation and each standard fully-connected layer is followed by batch normalization with momentum 0.9.
@@
Appendix C Embedding Space from the SAE applied on the $E$$t$ Maps
Figure 4 shows the embedding space obtained from the experimental SAE applied on the $E$$t$ maps.
Figure 4: Two-dimensional t-SNE projection of the learned latent space from the SAE applied on the $E$$t$ maps. Panel A: Points are color-coded by the variability index of the corresponding X-ray sources. Panel B: Points are color-coded by the hard-to-soft X-ray hardness ratio. Panel C: Known dips, flares, and pulsating sources (crosses) distributed across the embedding space.

### v2

@@ 2.1 Tensor Representation of Event Time Series
Assume that each event time series consists of a sequence of timestamps $\boldsymbol{t}=\{t_{k}\}_{k=1}^{N}$ and associated event modality values $\boldsymbol{E}=\{E_{k}\}_{k=1}^{N}$. In the context of X-ray astronomy, $\boldsymbol{t}$ would correspond to the photon arrival times and $\boldsymbol{E}$ to the photon energies. The total duration of the event time series is given by $T=t_{N}-t_{1}$, and the event values lie in the range $E_{\min},E_{\max}]$, where $E_{\min}=\min_{k}E_{k}$ and $E_{\max}=\max_{k}E_{k}$. To standardize the inputs across event time series, we normalize the time axis as
@@
These three normalized axes—event time ($\boldsymbol{\tau}$), transformed modality ($\boldsymbol{\epsilon}$), and inter-event interval ($\boldsymbol{\Delta\tau}$)—are discretized into fixed-width bins with $n_{\tau}$, $n_{\epsilon}$, and $n_{\Delta\tau}$ bins, respectively. Each event is assigned to a voxel in the resulting three-dimensional histogram based on its coordinates in this normalized space. The number of bins along each axis can be chosen flexibly to achieve the desired level of temporal or modality-specific resolution, depending on the application, domain, and characteristics of the data. This allows the representation to be tailored to the appropriate scale and granularity for a given task, while still providing a consistent input format for models, even when the dataset contains sequences of variable length and irregular temporal structure. This results in a tensor $\boldsymbol{X}_{i}\in\mathbb{R}^{n_{\tau}\times n_{\epsilon}\times n_{\Delta\tau}}$, which provides a standardized, structured input format for each of the original event time series $\boldsymbol{x}_{i}$. Each tensor is then collected into a final processed dataset $\{\boldsymbol{X}_{i}\}^{n}_{i=1}$. We also refer to these tensors as EEttdtdt cubes. If we were to only bin along time $\boldsymbol{\tau}$ and modality $\boldsymbol{\epsilon}$, we would obtain two-dimensional tensors $\boldsymbol{X}_{i}\in\mathbb{R}^{n_{\tau}\times n_{\epsilon}}$, which we refer to as EEtt maps. Appendix A shows EEtt maps and EEttdtdt cubes for an X-ray flare, dip and pulsating source.
Figure 1: Two-dimensional t-SNE projection of the learned latent space from the SAE applied on the EEtt-dtdt cubes. Panel A: Points are color-coded by the variability index of the corresponding X-ray sources. Panel B: Points are color-coded by the hard-to-soft X-ray hardness ratio. Panel C: Known dips, flares, and pulsating sources (crosses) occupy distinct clusters in the embedding space, enabling the identification of new transient candidates via clustering and similarity searches.
@@ 2.2 Sparse Representation Learning
We now aim to learn compact and informative representations from the EEttdtdt cubes. To this end, we employ a SAE—an unsupervised neural network model designed to extract structured, low-dimensional features from high-dimensional inputs. By learning to reconstruct the input data while enforcing sparsity in the latent space, the SAE identifies salient patterns and compresses each tensor into a representation that preserves essential temporal and modality-specific features, while remaining insensitive to irrelevant variation (Ranzato et al., 2007; Ng et al., 2011). This includes robustness to noise, systematics, observational artifacts, and context-dependent features introduced by the data acquisition process. Instead, the learned representations are encouraged to capture the most important latent factors that reflect the underlying physical characteristics and meaningful signals encoded in the time series, rather than superficial or incidental aspects of the data.
@@
where $\lambda>0$ controls the strength of the sparsity regularization. Once trained, we discard the decoder and use the encoder $\phi_{\theta}$ as a standalone feature extractor. The resulting latent vectors $\boldsymbol{z}_{i}$ serve as learned representations of the original event time series and can be directly used for downstream tasks such as clustering, anomaly detection, or similarity-based retrieval. A similar procedure is followed when using the EEtt maps as inputs to a SAE.
@@ 3 Experiments on Downstream Tasks
We apply our method to data from the Chandra Source Catalog (CSC) version 2.1 (Evans et al., 2024). Specifically, we use region event files, which list individual photon events for a specific source in a given observation—each with associated arrival time and energy. The dataset comprises 95,473 event files from 58,932 unique X-ray sources. For further details on the dataset, please refer to Dillmann et al. 2025. Using the terminology introduced in Section 2.1, the photon arrival times correspond to $\boldsymbol{t}$ and the photon energies to $\boldsymbol{E}$. We choose the event transformation function $f(\cdot)$ to be the base-10 logarithm, i.e., $f(\boldsymbol{E})=\log_{10}(\boldsymbol{E})$, which is commonly used in X-ray astronomy. For the EEttdtdt cubes, we choose a dimensionality of $(n_{\tau},n_{\epsilon},n_{\Delta\tau})=(24,16,16)$. For the SAE, we choose a sparsity strength of $\lambda=0.1$ and its latent vector dimension is 24. We also run the experiments described in this section on the EEtt maps with a dimensionality of $(n_{\tau},n_{\epsilon})=(24,16)$. In this case, we use a SAE with convolution layers and a latent vector dimension of 12. More details are provided in Appendix B.
@@
We train a vanilla XGBoost classifier and regressor (Chen & Guestrin, 2016) on the learned latent representations to distinguish highly variable X-ray sources (variability index >> 6) from non-variable ones (variability index $\leq$ 6), and to predict the hardness ratio of each source. The performances are summarized in Table 1. The classifier achieves a high variability classification accuracy of 0.97, outperforming the method proposed in Song et al. (2025), and the hardness ratio regressor attains a slightly reduced R2 score of 0.76, demonstrating that the learned representations encode temporal variability well and retain substantial predictive power for spectral hardness.
@@ 4 Conclusion
We introduce a unified framework for learning sparse representations from event time series to enable a variety of different downstream tasks. We represent irregular event time series as EEttdtdt cube tensors and employ a SAE to learn sparse embeddings that capture physically meaningful variation while increasing their robustness to noise and other irrelevant features. The method requires minimal manual tuning and scales to large datasets across various domains, and serves as a foundation for subsequent analysis like anomaly detection, similarity search, and unsupervised classification. Applied to archival Chandra X-ray data, the learned representations encode temporal and spectral information and enabled the discovery of a new extragalactic FXT and a TDE from a hyperluminous supersoft X-ray source, both previously missed by other transient search methods. Future work involves optimizing the resolution of the EEttdtdt cubes, the deployment of more expressive SAE architectures and the application of the method to a variety of other scientific datasets.
@@ Acknowledgements
Appendix A $E$-$t$ Maps and $E$-$t$-dtdt Cubes Visualizations
Figure 3 displays the light curves, EEtt maps, and EEttdtdt cubes for event files featuring a dip, a flare, and pulsations, respectively.
Figure 3: The top row shows 300 s bin light curves from event files featuring a dip (blue), a flare (red), and pulsations (green), respectively. The middle row shows the corresponding EEtt maps, and the bottom row shows the corresponding EEttdtdt cubes for these event files.
@@ Appendix B SAE Architecture and Training
The architecture of the experimental SAE for the EEttdtdt cube inputs is summarized in Table 2. For the experiments presented in this paper, we flatten the input and use fully-connected layers for computational reasons. We recommend experimenting with more sophisticated SAE architectures to fully leverage the value of the EEttdtdt cubes and the presented SAE representation learning approach. Starting with the original dataset of 95,473 samples, we first apply a 90/10 split to obtain 85,925 samples for training and validation, and 9,548 for testing. We then further split the training and validation subset using an 80/20 ratio, resulting in 68,740 training samples and 17,185 validation samples. The model is trained for up to 200 epochs with a batch size of 1024. Optimization is performed using Adam (Kingma & Ba, 2014) with an initial learning rate of $0.01$. A learning rate scheduler reduces the rate by a factor of 10 if the validation loss plateaus for more than 10 epochs. We also apply early stopping based on the validation loss, which halts training if no improvement is observed for 25 consecutive epochs and restores the best-performing model weights. A similar procedure is followed for the experimental SAE with convolution layers summarized in Table 3 that we used for the EEtt map inputs.
Table 2: Summary of the encoder architecture of the fully-connected autoencoder used to extract informative features from the $E$-$t$-dtdt cubes. Each layer uses Leaky ReLU activation and each standard fully-connected layer is followed by batch normalization with momentum 0.9.
@@
Appendix C Embedding Space from the SAE applied on the EEtt Maps
Figure 4 shows the embedding space obtained from the experimental SAE applied on the EEtt maps.
Figure 4: Two-dimensional t-SNE projection of the learned latent space from the SAE applied on the EEtt maps. Panel A: Points are color-coded by the variability index of the corresponding X-ray sources. Panel B: Points are color-coded by the hard-to-soft X-ray hardness ratio. Panel C: Known dips, flares, and pulsating sources (crosses) distributed across the embedding space.